A sum of money is divided among three people, X, Y and Z, in the ratio 5:3:1. If Y gets Shillings 700 more than Z, calculate how much X gets ?
Answer by ankor@dixie-net.com(22740)

A sum of money is divided among three people, X, Y and Z, in the ratio 5:3:1.
If Y gets Shillings 700 more than Z, calculate how much X gets?
:
let m be the multiplier
then each amt is:
x = 5m
y = 3m
z = m
:
"If Y gets Shillings 700 more than Z"
3m = m + 700
3m - m = 700
2m = 700
m = 350 is the multiplier
:
x = 5 * 350
x = 1750 shillings